# Just rephrase it! Uncertainty estimation in closed-source language models via multiple rephrased queries

### v1

@@ 1 Introduction
2 Rephrasings for top-1 decoding
@@
3 A model of the latent space
@@
We prove the above for the binary case of two classes $A$ and $B$ in Appendix C, but expect that it should be sufficiently informative in multi-class settings when $A,B$ are much more probable than other classes. A crucial assumption for recovering well-calibrated predictions is that $\boldsymbol{\mathrm{w}}^{\top}\epsilon_{rephrase}\sim\rho$ follows a logistic distribution with $\mu=0$ and $s=1$. We test this assumption by computing the cumulative of $\rho$ for our different experimental setups. In Figure 3(c) we find and plot the empirical cumulative using a Kolmogorov-Smirnov test (Smirnov, 1948) and $S=100$ MC samples of $\rho$ for Mistral-7B, ARC-Challenge, and the “expansion” rephrasing method. We see that the indeed the cumulative is approximately logistical validating our prediction (the confidence bands cover different queries $\boldsymbol{x}$).
4 Top-k decoding and temperature scaling
@@
The approximation relies on linearizing the involved functions, however, it is illustrative of the effect of both $s^{2}_{topk}$ and $s^{2}_{rephrase}$. In particular, we see that both $s^{2}_{topk}$ and $s^{2}_{rephrase}$ act to temper the probability $p(A|\boldsymbol{z}_{mean},f)$ under the categorical distribution of g. This highlights why using rephrasings with an appropriate temperature might improve the calibration in downstream tasks. In previous works, tempering of the categorical distribution has been found to significantly improve the calibration of deep neural networks (Guo et al., 2017). This also highlights a potential limitation of rephrasings in improving calibration. In particular, we expect that if we temper with an appropriate temperature the results obtained by top-k alone we should be able to match the best performance of rephrasings.
5 Experiments on top-k decoding
In Tables 5, 6 and 7 and Figure 3, we present the results for the top-k experiments with and without rephrasing, with $k=40$. We see that the stochasticity of top-40 compared to top-1 decoding from Tables 2, 3 and 4 results in an improvement in calibration but a drop in accuracy such that the Brier score often deteriorates. Further stochasticity in answers caused by rephrasings has a similar effect. These observations are consistent with the fact that top-k and nucleus sampling (Holtzman et al., 2019) make text more human-like but not necessarily more “accurate”. Overall, the Brier score in Tables 2, 3 4 and 5, 6 7 indicate that top-1 decoding with rephrasing is often the best in terms of a tradeoff between accuracy and calibration. However, if the main goal is calibration, the tables, and Figure 3 show that in accordance with proposition 4.1 rephrasing acts primarily to temper the probability of the top class. This often improves calibration significantly in terms of ECE, and AUROC especially for smaller models. Thus, when using top-k decoding and rephrasing, the user should assess whether rephrasing is appropriate after further analysis of his individual model and task.
Mistral-7B top-k 0.746, 0.272, 0.091, 0.511, 0.6, -
[…]
rephrase 0.64, 0.106, 0.086, 0.485, 0.82, 1.0
paraphrase 0.631, 0.11, 0.098, 0.495, 0.83, 1.0
expansion 0.517, 0.061, 0.114, 0.573, 0.859, 1.5
Llama-2-7B top-k 0.436, 0.201, 0.139, 0.761, 0.602, -
[…]
rephrase 0.356, 0.314, 0.17, 0.944, 0.627, 1.0
paraphrase 0.309, 0.185, 0.162, 0.851, 0.69, 1.5
expansion 0.322, 0.144, 0.155, 0.828, 0.622, 1.5
Llama-2-13B top-k 0.462, 0.125, 0.115, 0.679, 0.753, -
[…]
rephrase 0.398, 0.068, 0.136, 0.725, 0.743, 1.0
paraphrase 0.364, 0.109, 0.137, 0.738, 0.719, 1.2
expansion 0.373, 0.124, 0.143, 0.76, 0.669, 1.5
@@ 5 Experiments on top-k decoding
(a) $p_{A}(\boldsymbol{x})$ for top-k without rephrasing
(b) $p_{A}(\boldsymbol{x})$ for top-k with rephrasing
(c) Logistic (blue), and empirical cdf (red)

### v2

@@ 1 Introduction
2 Rephrasing drastically improves calibration for top-1 decoding
@@
In Appendix D, we also compare with Chain-of-Thought Wei et al. (2022) for uncertainty estimation. We find that we get competitive results with CoT. At the same time our method is significantly easier and more natural to implement for humans interacting via text with an LLM.
@@
Table 5: Comparisons between our rephrasing methods and white-box logit uncertainty estimation. We see that our rephrasing methods achieve similar calibration to what would be achieved if we had access to last layer logits. This is evident both in the AUROC and TACE as well as the Brier score, which also accounts for accuracy.
Dataset Model Method Acc $\uparrow$ ECE $\downarrow$ TACE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
ARC-C Mistral-7B logits 0.742 0.252 0.075 0.503 0.741
expansion 0.602 0.133 0.099 0.509 0.847
[…]
expansion 0.373 0.112 0.153 0.778 0.687
[…]
reword 0.445 0.084 0.119 0.714 0.721
ARC-E Mistral-7B logits 0.866 0.128 0.037 0.264 0.818
reword 0.753 0.045 0.062 0.297 0.931
[…]
rephrase 0.535 0.131 0.117 0.603 0.830
[…]
expansion 0.524 0.078 0.12 0.552 0.893
OBQA Mistral-7B logits 0.655 0.298 0.085 0.602 0.705
reword 0.552 0.105 0.102 0.592 0.796
[…]
expansion 0.362 0.083 0.138 0.775 0.678
[…]
rephrase 0.428 0.095 0.14 0.729 0.73
3 Rephrasing works as well as having access to the last layer logits
@@
We prove the above for the binary case of two classes $A$ and $B$ in Appendix C, but expect that it should be sufficiently informative in multi-class settings when $A,B$ are much more probable than other classes. A crucial assumption for recovering well-calibrated predictions is that $\boldsymbol{\mathrm{w}}^{\top}\epsilon_{rephrase}\sim\rho$ follows a logistic distribution with $\mu=0$ and $s=1$. We test this assumption by computing the cumulative of $\rho$ for our different experimental setups. In Figure 3(c) we find and plot the empirical cumulative using a Kolmogorov-Smirnov test (Smirnov, 1948) and $S=100$ MC samples of $\rho$ for Mistral-7B, ARC-Challenge, and the “expansion” rephrasing method. We see that the indeed the cumulative is approximately logistical validating our prediction (the confidence bands cover different queries $\boldsymbol{x}$). In Table 5 we use the logits of the answers as an oracle white-box uncertainty estimate. Specifically, we apply the softmax function and use the probability of the most probable class as our estimate of uncertainty. We compare the results of this method with the best rephrasing method (in terms of Brier) from Tables 2, 3 and 4. We see observe that our uncertainty estimates that are similar to what we would get if we had access to the last layer logits.
4 For top-k decoding, rephrasing tempers predictive uncertainty
(a) $p_{A}(\boldsymbol{x})$ for top-k without rephrasing
(b) $p_{A}(\boldsymbol{x})$ for top-k with rephrasing
(c) Logistic (blue), and empirical cdf (red)
Mistral-7B top-k 0.746, 0.272, 0.091, 0.511, 0.6, -
temp-sampling 0.742 0.272 0.089 0.513 0.605 -
[…]
rephrase 0.64, 0.106, 0.086, 0.485, 0.82, 1.0
paraphrase 0.631, 0.11, 0.098, 0.495, 0.83, 1.0
expansion 0.517, 0.061, 0.114, 0.573, 0.859, 1.5
Llama-2-7B top-k 0.436, 0.201, 0.139, 0.761, 0.602, -
temp-sampling 0.441 0.211 0.132 0.757 0.621 -
[…]
rephrase 0.356, 0.314, 0.17, 0.944, 0.627, 1.0
paraphrase 0.309, 0.185, 0.162, 0.851, 0.69, 1.5
expansion 0.322, 0.144, 0.155, 0.828, 0.622, 1.5
Llama-2-13B top-k 0.462, 0.125, 0.115, 0.679, 0.753, -
temp-sampling 0.47, 0.122 0.115 0.662 0.766 -
[…]
rephrase 0.398, 0.068, 0.136, 0.725, 0.743, 1.0
paraphrase 0.364, 0.109, 0.137, 0.738, 0.719, 1.2
expansion 0.373, 0.124, 0.143, 0.76, 0.669, 1.5
@@ 4 For top-k decoding, rephrasing tempers predictive uncertainty
The approximation relies on linearizing the involved functions, however, it is illustrative of the effect of both $s^{2}_{topk}$ and $s^{2}_{rephrase}$. In particular, we see that both $s^{2}_{topk}$ and $s^{2}_{rephrase}$ act to temper the probability $p(A|\boldsymbol{z}_{mean},f)$ under the categorical distribution of g. This highlights why using rephrasings with an appropriate temperature might improve the calibration in downstream tasks. In previous works, tempering of the categorical distribution has been found to significantly improve the calibration of deep neural networks (Guo et al., 2017).
In Tables 6, 7 and 8 and Figure 3, we present the results for the top-k experiments with and without rephrasing, with $k=40$. We also present the relaxed temperature sampling variant Wei et al. (2022). We see that the stochasticity of top-40 compared to top-1 decoding from Tables 2, 3 and 4 results in an improvement in calibration but a drop in accuracy. The Brier score often improves at the cost of accuracy. Further stochasticity in answers caused by rephrasings has a similar effect. These observations are consistent with the fact that top-k and nucleus sampling (Holtzman et al., 2019) make text more human-like but not necessarily more “accurate”. However, if the main goal is calibration, the tables, and Figure 3 show that in accordance with proposition 4.1 rephrasing acts primarily to temper the probability of the top class. This often improves calibration significantly in terms of ECE, and AUROC especially for smaller models. We plot the results of all methods averaged over all models for each dataset in Figure 4. Tables 2, 3 4 and 6, 7 8 and Figure 4 indicate that the user should assess whether rephrasing is appropriate after an analysis of his individual model, task and evaluation metric. However, in general, a hyperparameter-optimized choice of rephrasing + top-1 decoding outperforms or matches all other method combinations in all metrics.
@@ 5 Related works
(a) ARC-Challenge
(b) ARC-Easy
(c) OpenBookQA
Figure 4: We plot the AUROC averaged over all models for each dataset and for each uncertainty estimation method. We observe that top-k improves over the naive top-1 decoding. Furthermore, the best rephrasing method (denoted as rephrase*) improves the AUROC significantly in all cases.
@@ Appendix C Additional Proofs
Appendix D Additional comparisons with CoT
We compare with Chain-of-Thought Wei et al. (2022) for uncertainty estimation and plot the results in Table 13. We find that we get competitive results with CoT. At the same time our method is significantly easier and more natural to implement for humans interacting via text with an LLM. In CoT one needs to first obtain a sequence of reasoning steps. These should then be used as additional context when asking an LLM to answer again the base question. By contrast we propose a simple one step process of rephrasing the base question.
Table 13: Comparisons between our best rephrasing method and CoT. Our rephrasing method obtains comparable results to CoT in terms of Brier score and other calibration metrics.
Dataset Model Method Acc $\uparrow$ ECE $\downarrow$ TACE $\downarrow$ Brier $\downarrow$ AUROC $\uparrow$
ARC-C Mistral-7B CoT 0.725 0.173 0.071 0.439 0.719
expansion 0.602 0.133 0.099 0.509 0.847
[…]
expansion 0.373 0.112 0.153 0.778 0.687
[…]
reword 0.445 0.084 0.119 0.714 0.721
ARC-E Mistral-7B CoT 0.857 0.07 0.037 0.211 0.829
reword 0.753 0.045 0.062 0.297 0.931
[…]
rephrase 0.535 0.131 0.117 0.603 0.830
[…]
expansion 0.524 0.078 0.12 0.552 0.893
OBQA Mistral-7B CoT 0.662 0.153 0.083 0.501 0.762
reword 0.552 0.105 0.102 0.592 0.796
[…]
expansion 0.362 0.083 0.138 0.775 0.678
[…]
rephrase 0.428 0.095 0.14 0.729 0.73